OpenStudy (anonymous):

solve the equation w=Cr^-2 for r

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

go to myalgebra.com

OpenStudy (anonymous):

i dont know how to write the answer to u

OpenStudy (anonymous):

ok

OpenStudy (jamesj):

(Rhaukus, could you let me know when you have an answer to your Gamma function question? I'm intrigued.)

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

\[w=Cr^{-2}\]\[w=C/r^{2}\]\[wr^2=C\]\[r^2 = C/w\]\[r=\sqrt(C/w)\]

OpenStudy (unklerhaukus):

@ jamesj im still working on that Gamma function question, i will let you know when i have solved it

OpenStudy (anonymous):

this is what i had is it the same r=√wC/W

OpenStudy (anonymous):

0K

OpenStudy (jamesj):

(@rh, thanks.)